If a, b and c are the roots of the equation x^3 -6x^2 +10x-6 =0 , find the values of a^2 +b^2 +c^2 and 1/a + 1/b + 1/c.


1
Expert's answer
2021-11-16T11:08:22-0500

If "a, b" and "c" are the roots of the equation "x^3 -6x^2 +10x-6 =0 ," let us find the values of "a^2 +b^2 +c^2" and "\frac{1}a + \frac{1}b + \frac{1}c."

According to Vieta's formulas, "a+b+c=6,\ ab+ac+bc=10,\ abc=6."

Then we get

"a^2 +b^2 +c^2=(a+b+c)^2-2(ab+bc+ac)=6^2-2\cdot 10=36-20=16"

and

"\frac{1}a + \frac{1}b + \frac{1}c=\frac{bc+ac+ab}{abc}=\frac{10}{6}=\frac{5}3."